LO10-2.L010-2, PROBLEM 108-3 Comprehensive Variance Analysis: Journal Entries [LO10-1, L01 Wandard cost vyv alls. The standar i was just wondering if you could completely walk through this problem for me thanks
Answer #1

Part 1 A

Materials price variance= AQ*(AP-SP) = 32000*(4.80-5) = -6400 =$6400 F

Materials quantity variance= SP*(AQ-SQ) = 5.00*(32000-29600) = $12000 U

Part 1 B

No. Account titles and explanation debit credit
1 raw materials (32000*5) 160000
Materials price variance 6400
Accounts payable (32000*4.80) 153600
2 work in process (29600*5) 148000
Materials quantity variance 12000
Raw materials (32000*5) 160000

Part 2 a

Labor rate variance= AH*(AR-SR) = 6400*(8-7.50) = 3200 U

Labor efficiency variance= SR*(AH-SH) = 7.50*(6400-7200) = -6000 = $6000 F

Part 2 B

No. Account titles and explanation Debit credit
1 work in process (7200*7.50) 54000
Labor rate variance 3200
Labor efficiency variance 6000
Wages payable (6400*8) 51200

Part 3

Variable overhead spending variance= AH*(AR-SR) = 6400*(2.75-2.50) = 1600 U

Variable overhead efficiency variance= SR*(AH-SH) = 2.50*(6400-7200) = -2000= $2000 F

Part 4

His statement is not correct because there is large unfavourable material quantity variance and overhead spending variance. The idea to increase output by giving raises to employees is a good idea but its implementation didn't prove to be effective as it resulted into high labor rate variance

Part 5

Possible reasons for variance

favourable material price variance: inferior, quality materials drop in market price, discount on quantity purchased

Unfavourable materials quantity variance: inferior quality materials and skilled workers and carelessness

Unfavourable labour rate variance: use of highly skilled workers, pay scale changes, overtime working hours

Favourable labour efficiency variance: high quality of materials, highly skilled employees, use of properly maintained equipments

Unfavourable variable overhead spending variance: trees in wastage, uneconomical purchase, increase in cost

Favourable variable overhead efficiency variance: high quality of materials, highly Skilled employees, use of properly maintained equipments